suppose the coordinate of p=2 and PQ=8 and PR=12. Whare are the possible mid point for line segment PQ?
1
Expert's answer
2013-08-06T08:12:07-0400
Let the coordinate of Q is x.
We know that coordinate of P is 2.
Then the coordinate of midpoint is 0.5( P+Q)=0.5( 2 + x)=1+0.5x On other hand PQ=8 so | x - 2|=8 there are 2 cases x-2=8 or x-2=-8 so
x=10 or x=-6 Then coordinate of midpoint equals 1+0.5x .For x=10 we get 6 andfor x=-6 we get -2
Answer: coordinate of midpoint of PQ is 6 or -2